A particle has a net charge of +0.31 nC.

How many electrons must be added to this particle in order to make it electrically neutral? (to 2 s.f)

[NOTE: enter numbers using scientific notation by using "E": for example 1.6×10−19 would be entered as 1.6E-19]
1
Expert's answer
2020-05-19T10:59:28-0400

We should add the total charge equal to -0.31nC or

"-0.31\cdot E{-9}\, \mathrm{C}." This charge should be equal to "N\cdot e," where "N" is the number of electrons and "e" is the charge of one electron. So the number will be

"N = \dfrac{-0.31E-9\,\mathrm{C}}{e} = \dfrac{-3.1E-10\,\mathrm{C}}{-1.6E-19\,\mathrm{C}} \approx 1.94\cdot10^9."